Your current income is equal to 60,00. Your next period (future) income is known to be equal to 48,000. If your current consumption expenditure is equal to 48,000, what is your (current) level of savings? Ok I got this answer, which is S=12,000. However I cannot get the following one, which is: If the real rate of interest is equal to 10%, how much will you spend on consumption next period (assuming that your current consumption is 48,000)?
Answer #1

Savings accumulated at 10% = 12000*1.10 = 13200

Thus, spenpding on next period consumption=

Accumulated savings + next period income

= 13200+ 48000 = 61200.

Thanks!